Sample with size n = 100 has mean = 30. Assuming the population standard deviation is 8, construct 95% confidence interval for population mean. Use formula zα/2 = 1.96
Answer ? Help?

between 28.4 and 31.6

between 27.0 and 33.0

between 26.1 and 34.2

between 24.1 and 35.9


Answer by jim_thompson5910(35256)

Lower Limit:

L = mean - za/2 * (standard deviation)/sqrt(n)

L = 30 - 1.96 * (8)/sqrt(100)

L = 30 - 1.96 * (8)/10

L = 30 - 1.96 * 0.8

L = 30 - 1.568

L = 28.432


Upper Limit:

U = mean + za/2 * (standard deviation)/sqrt(n)

U = 30 + 1.96 * (8)/sqrt(100)

U = 30 + 1.96 * (8)/10

U = 30 + 1.96 * 0.8

U = 30 + 1.568

U = 31.568



So the 95% confidence interval is (L, U) = (28.432, 31.568)

Round this to one decimal place to get (28.4, 31.6)

So the answer is Choice A) between 28.4 and 31.6